It is known that quantity demanded decreases by two units for each $1 increase in price. At a price of $5, quantity demanded is ten units.
a) What will be the quantity demanded is price is zero?
b) Write and equation for quantity demanded as a function of price.
c) Write an equation that expresses price as a function of quantity.
d) Write an equation for total revenue.
1
Expert's answer
2015-07-23T00:00:42-0400
It is known that quantity demanded decreases by two units for each $1 increase in price. At a price of $5, quantity demanded is ten units.
a) If the price is zero the quantity demanded will be zero too.
b) The equation for quantity demanded as a function of price is Q = 20 - 2P.
c) An equation that expresses price as a function of quantity will be P = 10 - 0.5Q.
d) An equation for total revenue is:
TR = P*Q = (10 - 0.5Q)*Q = 10Q - 0.5Q^2.